# Preserving the Hermiticity of the One-Body Density Matrix for a Non-Interacting Fermi Gas

### v2

@@ 4 Concluding remarks
In summary we have studied the approximate semiclassical ODM of a zero temperature non-interacting Fermi gas obtained through three different gradient expansion techniques: the KODM, WKODM, and GVODM methods. We have resolved the inconsistencies discussed in the literature [1, 21, 22], and have shown that, regardless of whichever method is used, the ODMs are Hermitian, idempotent, and identical when cast into symmetric coordinates $(\mathbf{R},\mathbf{s})$. Inconsistencies only arise when the semiclassical expansions are truncated in terms of non-symmetric coordinates $(\mathbf{r},\mathbf{r}^{\prime})$, which occur with the non-symmetric KODM [equation 2.1] and WKODM methods [A]. Additionally, we have provided two separate and explicit derivations of the $d\geq 1$-dimensional GVODM originally proposed by reference [1] based on their $d=1,2,3,4$ expressions. The first derivation (Section 2) was accomplished by transforming the known $d\geq 1$-dimensional expression of the KODM [20] into symmetric coordinates, while the second derivation (Section 3) transformed the known $d\geq 1$-dimensional WKODM into symmetric coordinates from the very beginning. For simplicity, we only dealt with semiclassical terms up to second order in $\hbar$, nevertheless, our results still hold when higher order terms in the ODMs are accounted for.
@@
Appendix A Non-symmetric WKODM
In this section, we apply the WKODM method in non-symmetric coordinates, and obtain an expression identical to the non-symmetric KODM of equation 2.1, finding that $\rho_{WK}(\mathbf{r},\mathbf{r}^{\prime})=\rho_{K}(\mathbf{r},\mathbf{r}^{\prime})$. This is not to be confused with our symmetric WKODM method derivation in Section 3, where we similarly show that $\rho_{WK}(\mathbf{R},\mathbf{s})=\rho_{GV}(\mathbf{R},\mathbf{s})$.
In showing $\rho_{WK}(\mathbf{r},\mathbf{r}^{\prime})=\rho_{K}(\mathbf{r},\mathbf{r}^{\prime})$, we see that equations 2.2, 2.3 and 2.4 can alternatively be expressed in terms of inverse Laplace transformations via equations 2.5, 2.6 and 2.7. Using this fact we are able to show in Section 2 how to properly transform the KODM into symmetric coordinates $(\mathbf{R},\mathbf{s})$ by performing the transformation of variables before the inverse Laplace transforms are taken. Making the transformation after would instead result in additional terms being incorrectly unaccounted for.
@@ Appendix A Non-symmetric WKODM
In the non-symmetric WKODM method the $d\geq 1$-dimensional Bloch matrix is approximated to second order in $\hbar$ by $C(\mathbf{r},\mathbf{r}^{\prime};\beta)\approx C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)+C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)+C_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ where
@@
$C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=\frac{g}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}},$ (1.1)
$C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=-\frac{ig\hbar\beta^{2}}{2m(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}w_{1}(\mathbf{r},\mathbf{p};\beta),$ (1.2)
$C_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=\frac{g\hbar^{2}}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}w_{2}(\mathbf{r},\mathbf{p};\beta),$ (1.3)
and
$w_{1}(\mathbf{r},\mathbf{p};\beta)$ $=-\bigg{(}\frac{i\beta^{2}}{2m}\bigg{)}\big{(}\mathbf{p}\cdot\nabla_{\mathbf{r}}V(\mathbf{r})\big{)},$ (1.4)
$w_{2}(\mathbf{r},\mathbf{p};\beta)$ $=-\frac{\beta^{3}}{4m}\nabla_{\mathbf{r}}^{2}V(\mathbf{r})+\frac{\beta^{3}}{6m}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\big{)}^{2}$
$\quad+\frac{\beta^{3}}{6m^{2}}(\mathbf{p}\cdot\nabla_{\mathbf{r}})^{2}V(\mathbf{r})-\frac{\beta^{4}}{8m^{2}}\big{(}\mathbf{p}\cdot\nabla_{\mathbf{r}}V(\mathbf{r})\big{)}^{2}.$ (1.5)
$H_{cl}(\mathbf{r},\mathbf{p})=\mathbf{p}^{2}/(2m)+V(\mathbf{r})$ and the details of where these expressions come from are described in reference [10], although they focus on the specific case of a diagonal system where $\mathbf{r}=\mathbf{r}^{\prime}$ so that $\mathbf{s}=0$. Our case is more general since we assume a small but finite $\mathbf{s}$. Our symmetric WKODM derivation in Section 3 will also result in similar expressions, but in terms of symmetric coordinates.
Combining equations 1.1, 1.2, 1.3, 1.4 and 1.5 is precisely the origin of equations 2.8, 2.9 and 2.10. In order to justify expressing the KODM in terms of inverse Laplace transforms we show below that in non-symmetric coordinates the KODM [equations 2.1, 2.2, 2.3 and 2.4] is exactly obtained by substituting equations 2.8, 2.9 and 2.10 into equations 2.5, 2.6 and 2.7, respectively. Starting with the $A_{0}$ term, equations 2.5 and 2.8 yield
$A_{0}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}}$ $=g\bigg{(}\frac{m}{2\pi\hbar^{2}}\bigg{)}^{d/2}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2-1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
$=\frac{gk_{F}^{d}}{(2\pi z)^{d/2}}J_{\frac{d}{2}}(z).$
Similarly for $A_{1}$, equations 2.6 and 2.9 give
@@
$A_{1}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}}$ $=-\frac{gm^{d/2-1}\hbar^{2-d}}{2(2\pi)^{d/2}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\nabla_{\mathbf{s}}\big{)}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
$=-\frac{gk_{F}^{d-3}z^{2}}{4(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{(}\nabla_{\mathbf{r}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}.$
For $A_{2}$, combing equations 2.7 and 2.10 returns
$A_{2}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}}$ $=-\frac{gm^{d/2-1}}{12(2\pi)^{d/2}\hbar^{d-2}}\big{(}\nabla_{\mathbf{r}}^{2}V(\mathbf{r})\big{)}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
$\quad+\frac{gm^{d/2-1}}{24(2\pi)^{d/2}\hbar^{d-2}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\big{)}^{2}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+2}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
@@
$\quad+\frac{gm^{d/2}}{8(2\pi)^{d/2}\hbar^{d}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot{\mathbf{s}}\big{)}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
@@
$\quad-\frac{gm^{d/2}}{6(2\pi)^{d/2}\hbar^{d}}\Big{(}\nabla_{\mathbf{r}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot{\mathbf{s}}\big{)}\cdot{\mathbf{s}}\Big{)}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
$\implies A_{2}$ $=\frac{gk_{F}^{d-4}z^{2}}{24(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\big{(}\nabla_{\mathbf{r}}^{2}k_{F}^{2}\big{)}+\frac{gk_{F}^{d-6}z^{3}}{96(2\pi z)^{d/2}}J_{\frac{d}{2}-3}(z)\big{(}\nabla_{\mathbf{r}}k_{F}^{2}\big{)}^{2}$
@@
$\quad+\frac{gk_{F}^{d-6}z^{4}}{32(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\bigg{(}\nabla_{\mathbf{r}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}^{2}+\frac{gk_{F}^{d-4}z^{3}}{12(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{[}\nabla_{\mathbf{r}}\bigg{(}\nabla_{\mathbf{r}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}\bigg{]}.$
Each term was computed in terms of non-symmetric coordinates so that $k_{F}=k_{F}(\mathbf{r})$, and the inverse Laplace transform identity equation 2.4 was used in each case (with $\mathbf{R}$ replaced by $\mathbf{r}$). Additionally for the $A_{2}$ term, two $d\geq 1$-dimensional Gaussian integrals similar to that in equation 2.8 were solved straightforwardly, and lastly the vector identity $(\mathbf{s}\cdot\nabla_{\mathbf{r}})^{2}V(\mathbf{r})=\nabla_{\mathbf{r}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\mathbf{s}\big{)}\cdot\mathbf{s}$ was applied.

### v3

@@ 3 Concluding remarks
In summary we have studied the approximate semiclassical ODM of a zero temperature non-interacting Fermi gas obtained through three different gradient expansion techniques: the KODM, WKODM, and GVODM methods. We have resolved the inconsistencies discussed in the literature [23, 24, 25], and have shown that, regardless of whichever method is used, the ODMs are Hermitian, idempotent, and identical when cast into symmetric coordinates $(\mathbf{R},\mathbf{s})$. Inconsistencies only arise when the semiclassical expansions are truncated in terms of non-symmetric coordinates $(\mathbf{r},\mathbf{r}^{\prime})$, which occur with the non-symmetric KODM [equation 2.1] and WKODM methods. Additionally, we have provided an explicit derivation of the $d$-dimensional GVODM originally proposed by reference [23] based on their $d=1,2,3,4$ expressions. This was accomplished in Section 2 by showing that the KODM and WKODM are equivalent in non-symmetric coordinates, and proceeding to transform the known $d$-dimensional expression of the WKODM into symmetric coordinates. For completeness, in C we also provided an alternate yet equivalent derivation, starting from the definition of the Wigner-Kirkwood Bloch density matrix. For simplicity, we only dealt with semiclassical terms up to second order in $\hbar$, nevertheless, our results still hold when higher order terms in the ODMs are accounted for.
@@
Author Contributions Statement
B.V.Z. and K.B. conceived the research, and performed the initial calculations. L.M.F., P.C., and D.E. completed the calculations, and wrote the manuscript.
Competing Interests Statement
The authors have no competing interests to declare.
Appendix A Equivalence of KODM and WKODM
The aim of this appendix is to apply the Wigner-Kirkwood expansion of the Bloch density matrix to derive, in non-symmetric coordinates, an expression of the WKODM and to prove the equivalence of the WKODM and KODM in $d$-dimensions, $\rho_{WK}(\mathbf{r},\mathbf{r}^{\prime})=\rho_{K}(\mathbf{r},\mathbf{r}^{\prime})$, given explicitly by equation 2.1.
@@ Appendix A Equivalence of KODM and WKODM
In the non-symmetric coordinates the $d$-dimensional Bloch matrix is approximated to second order in $\hbar$ by $C^{WK}(\mathbf{r},\mathbf{r}^{\prime};\beta)\approx C_{0}^{WK}(\mathbf{r},\mathbf{r}^{\prime};\beta)+C_{1}^{WK}(\mathbf{r},\mathbf{r}^{\prime};\beta)+C_{2}^{WK}(\mathbf{r},\mathbf{r}^{\prime};\beta)$, and as shown in Section 2 the resulting WKODM is written as $\rho_{WK}(\mathbf{r},\mathbf{r}^{\prime})=A_{0}^{WK}+A_{1}^{WK}+A_{2}^{WK}$, where the $A_{0,1,2}^{WK}$ are respectively given equations 2.6, 2.7 and 2.8 together with equations 2.9, 2.10 and 2.11. We now proceed with the evaluation of each $A_{0,1,2}^{WK}$ term. Using equation 2.6 with equation 2.9 we have
$A_{0}^{WK}=g\bigg{(}\frac{m}{2\pi\hbar^{2}}\bigg{)}^{d/2}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2-1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}},$
and making use of the identity of equation 2.4 for $\xi=1$, we can write
$A_{0}^{WK}=\frac{gk_{F}^{d}}{(2\pi z)^{d/2}}J_{\frac{d}{2}}(z),$
with $z=k_{F}s$ and $k_{F}=\frac{\sqrt{2m}}{\hbar}(\mu-V(\mathbf{r}))^{1/2}$. This above result is identical to that given in equation 2.2 for the first term $A_{0}^{K}$ of the KODM, so that
$A_{0}^{WK}=A_{0}^{K}.$
Let us now come to the second term $A_{1}^{WK}$, given by equation 2.7 with equation 2.10. Using the identity $\nabla_{\mathbf{s}}\textrm{e}^{i\mathbf{p}\cdot\mathbf{s}/\hbar}=\frac{i\mathbf{p}}{\hbar}\textrm{e}^{i\mathbf{p}\cdot\mathbf{s}/\hbar}$, we can transform equation 2.10 as follows
@@
$C_{1}^{WK}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=-\frac{g\hbar^{2}\beta^{2}}{2m(2\pi\hbar)^{d}}\Big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\nabla_{\mathbf{s}}\Big{)}\int d^{d}p\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}$
$=-\frac{\hbar^{2}\beta^{2}}{2m}\Big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\nabla_{\mathbf{s}}\Big{)}C_{0}^{WK}(\mathbf{r},\mathbf{r}^{\prime};\beta)$
@@
$=-\frac{g\hbar^{2}\beta^{2}}{2m}\bigg{(}\frac{m}{2\pi\beta\hbar^{2}}\bigg{)}^{d/2}\Big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\nabla_{\mathbf{s}}\Big{)}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}$
@@
$=\frac{g\beta}{2}\bigg{(}\frac{m}{2\pi\beta\hbar^{2}}\bigg{)}^{d/2}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\mathbf{s}\big{)}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})},$
where we have successively used equation 2.9 and equation 2.12. Combining equation 2.7 with the above result, we can write
@@
$A_{1}^{WK}={\frac{g}{2}\bigg{(}\frac{m}{2\pi\hbar^{2}}\bigg{)}^{d/2}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot{\mathbf{s}}\big{)}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{{-d/2}}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}},$
applying once again the identity of equation 2.4, this time for $\xi=0$, we find
$A_{1}^{WK}=\frac{g}{2}\bigg{(}\frac{m}{2\pi\hbar^{2}}\bigg{)}^{d/2}\bigg{(}\frac{\hbar^{2}k_{F}}{ms}\bigg{)}^{d/2-1}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\mathbf{s}\big{)}J_{\frac{d}{2}-1}(z).$
By using the relation $\nabla_{\mathbf{r}}V(\mathbf{r})=-\frac{\hbar^{2}}{2m}\nabla_{\mathbf{r}}k_{F}^{2}$ and recalling that $z=k_{F}s$, we end with
$A_{1}^{{WK}}=-\frac{gk_{F}^{d-3}z^{2}}{4(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{(}\nabla_{\mathbf{r}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}.$
This expression is identical to that given in equation 2.3 for the second term $A_{1}^{K}$ of the KODM, thus
$A_{1}^{WK}=A_{1}^{K}.$
In a similar fashion we now calculate $A_{2}^{WK}$ of equation 2.7. We first rewrite equation 2.11 as
$C_{2}^{WK}(\mathbf{r},\mathbf{r}^{\prime};\beta)=$ $\bigg{(}-\frac{\beta^{2}}{4m}\big{(}\nabla_{\mathbf{r}}^{2}V(\mathbf{r})\big{)}+\frac{\beta^{3}}{6m}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\big{)}^{2}\bigg{)}C_{0}^{WK}(\mathbf{r},\mathbf{r}^{\prime};\beta)$
$+\frac{\hbar^{2}\beta^{4}}{8m^{2}}\Big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\nabla_{\mathbf{s}}\Big{)}^{2}C_{0}^{WK}(\mathbf{r},\mathbf{r}^{\prime};\beta)$
$-\frac{\hbar^{4}\beta^{3}}{8m^{2}}\Big{(}\nabla_{\mathbf{r}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\nabla_{\mathbf{s}}\big{)}\cdot\nabla_{\mathbf{s}}\Big{)}C_{0}^{WK}(\mathbf{r},\mathbf{r}^{\prime};\beta).$
We substitute the above expression into equation 2.7 together with equation 2.11 and obtain
@@
$A_{2}^{WK}=$ $-{\frac{g\hbar^{2}}{12m}\bigg{(}\frac{m}{2\pi\hbar}\bigg{)}^{d/2}}\big{(}\nabla_{\mathbf{r}}^{2}V(\mathbf{r})\big{)}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
$+{\frac{g\hbar^{2}}{24m}\bigg{(}\frac{m}{2\pi\hbar}\bigg{)}^{d/2}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\big{)}^{2}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+2}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
$+{\frac{g}{8}\bigg{(}\frac{m}{2\pi\hbar}\bigg{)}^{d/2}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot{\mathbf{s}}\big{)}^{2}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2+1}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}}$
$-{\frac{g}{6}\bigg{(}\frac{m}{2\pi\hbar}\bigg{)}^{d/2}}\Big{(}\nabla_{\mathbf{r}}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot{\mathbf{s}}\big{)}\cdot{\mathbf{s}}\Big{)}\mathcal{L}_{\mu}^{-1}\bigg{\{}\beta^{-d/2}\textrm{e}^{-\frac{m\mathbf{s}^{2}}{2\hbar^{2}\beta}-\beta V(\mathbf{r})}\bigg{\}},$
where we again have used the identity $\nabla_{\mathbf{s}}\textrm{e}^{i\mathbf{p}\cdot\mathbf{s}/\hbar}=\frac{i\mathbf{p}}{\hbar}\textrm{e}^{i\mathbf{p}\cdot\mathbf{s}/\hbar}$. To evaluate the various inverse Laplace transforms above, we will return to the identity of equation 2.4 using $\xi=0,-1$, and $-2$ this time, and find
$A_{2}^{WK}=$ $-\frac{g\hbar^{2}}{12m}\bigg{(}\frac{m}{2\pi\hbar^{2}}\bigg{)}\big{(}\nabla_{\mathbf{r}}^{2}V(\mathbf{r})\big{)}\bigg{(}\frac{\hbar^{2}k_{F}}{ms}\bigg{)}^{d/2-2}J_{\frac{d}{2}-2}(z)$
$+\frac{g\hbar^{2}}{24m}\bigg{(}\frac{m}{2\pi\hbar^{2}}\bigg{)}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\big{)}^{2}\bigg{(}\frac{\hbar^{2}k_{F}}{ms}\bigg{)}^{d/2-3}J_{\frac{d}{2}-3}(z)$
$+\frac{g}{8}\bigg{(}\frac{m}{2\pi\hbar^{2}}\bigg{)}\bigg{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\frac{\mathbf{s}}{s}\bigg{)}^{2}\bigg{(}\frac{\hbar^{2}k_{F}}{ms}\bigg{)}^{d/2-2}J_{\frac{d}{2}-2}(z)$
$-\frac{g}{6}\bigg{(}\frac{m}{2\pi\hbar^{2}}\bigg{)}\bigg{[}\nabla_{\mathbf{r}}\bigg{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\mathbf{s}\bigg{)}\cdot\mathbf{s}\bigg{]}\bigg{(}\frac{\hbar^{2}k_{F}}{ms}\bigg{)}^{d/2-1}J_{\frac{d}{2}-1}(z).$
Similarly to before, eliminating the potential $V(\mathbf{r})$ in favor of $k_{F}^{2}$ through the use of $\nabla_{\mathbf{r}}V(\mathbf{r})=-\frac{\hbar^{2}}{2m}\nabla_{\mathbf{r}}k_{F}^{2}$ allows us to arrange the above in the final form
$A_{2}^{WK}$ $=\frac{gk_{F}^{d-4}z^{2}}{24(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\big{(}\nabla_{\mathbf{r}}^{2}k_{F}^{2}\big{)}+\frac{gk_{F}^{d-6}z^{3}}{96(2\pi z)^{d/2}}J_{\frac{d}{2}-3}(z)\big{(}\nabla_{\mathbf{r}}k_{F}^{2}\big{)}^{2}$
$\quad+\frac{gk_{F}^{d-6}z^{4}}{32(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\bigg{(}\nabla_{\mathbf{r}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}^{2}+\frac{gk_{F}^{d-4}z^{3}}{12(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{[}\nabla_{\mathbf{r}}\bigg{(}\nabla_{\mathbf{r}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)}\cdot\frac{\mathbf{s}}{s}\bigg{]},$
an expression identical to equation 2.4, which allows us to write
$A_{2}^{WK}=A_{2}^{K}.$
Finally, combining the above results for $A_{0,1,2}^{WK}$, we have proved the equivalence
$\rho_{WK}(\mathbf{r},\mathbf{r}^{\prime})=\rho_{K}(\mathbf{r},\mathbf{r}^{\prime}).$